OpenStudy (firejay5):

Solve each system of equations by using substitution: Show work and step! Please! :D I worked the top equation and I am to the point where I am entering into the bottom equation. <--- I need help on that part. 16. - x - y = -2 4x + 5y = 16

OpenStudy (compassionate):

http://www.purplemath.com/modules/systlin4.htm

OpenStudy (firejay5):

@Compassionate sorry won't help me

OpenStudy (anonymous):

-y=-2+x all divided by -1 so then y=2-x then plug into other one: 4x + 5(2-x) =16 4x + 10-5x =16 subtract 10 4x-5x=6 -x=16 x=-16

OpenStudy (firejay5):

combine like terms ----> 4x - 5x

OpenStudy (firejay5):

where did you get - 16

OpenStudy (anonymous):

yes you combine the like terms

OpenStudy (anonymous):

from the previous process x=-16

OpenStudy (firejay5):

@hardequalsmath would x = -6 or x = 6

OpenStudy (anonymous):

oh shoot your right! I meant -6 Whoops! so then you plug -6 in.

OpenStudy (anonymous):

SO:-6-y=-2 -y=4 y=-4

OpenStudy (firejay5):

wouldn't you plug in -6 to -x make it a positive

OpenStudy (anonymous):

oh yeah sorry agian yes I didnt see the negative so 6-y=-2 -y=-8 y=8

OpenStudy (firejay5):

-y = 8 y = -8 ----> divided by -1

OpenStudy (firejay5):

@hardequalsmath sorry man!